An object having a velocity 4.0 m/s is accelerated at the rate of 1.2 m/s2 for 5.0 s. Find the distance travelled during the period of acceleration.

Grade:11

2 Answers

Aditi Chauhan
9 years ago
Sol. u = 4 m/s, a = 1.2 m/s2, t = 5 sec Distance = s = ut = + ½ at2 = 4(5) + 1/2 (1.2)52 = 35 m.
Rizwan
5 years ago
Sol. Given u=4m/s , a = 1.2 m/sand time = 5 secs. 
Distance = s = ut + 1/at2 = 4(5)+1/(1.2)(5)2
                            =20 + 1/2  (1.2) (25) 
                       = 20 + 3 (5) 
                        = 35 m 
                                Ans.